OpenStudy (anonymous):

Solve: y = x2 + 8x + 9 x - y = -3

OpenStudy (anonymous):

is it y=x^2 + 8x +9 ?

hero (hero):

It usually is

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

transpose the variable y in the equation x-y=-3

OpenStudy (anonymous):

put \[y=x+3\] in first equation to get \[x+3=x^2+8x+9\] hen solve the quadratic

OpenStudy (anonymous):

you solve \[x^2+7x+6=0\] \[(x+6)(x+1)=0\] \[x=-6,x=-1\]